# A Resolution Independent Neural Operator

### v2

@@ 5 Acknowledgements
We acknowledge the discussion with Professor George E. Karniadakis during the CRUNCH seminar, who brought random projection and ReNO [3] to our attention, which improved our work. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number DE-SC0024162.
@@
Complementary results for the numerical examples are provided in Appendices A, B and C. Appendix E lists the hyperparameters used for each problem. An ablation study for Algorithm 1, comparing our method with GPOD, is presented in Appendix F. Moreover, the use of random projections as an alternative for finding vector representations of input or output function data has been studied and compared to our proposed dictionary learning algorithm. Appendix G summarizes the formulation for Gaussian random field generation.
@@ Appendix A Nonlinear 1D Darcy’s Equation
The history of the loss function during the operator learning stage is plotted in Figure 15(a). In Figure 15(b), the distribution of output function prediction errors after training is shown. The results suggest that the training is successful and the output function prediction via RINO is very accurate.
@@ Appendix B Nonlinear 2D Darcy’s Equation
Two examples of the test predictions for the output field are shown in Figures 18 and 19, corresponding to the 75th percentile error and the median error, respectively. In each example, the queried point cloud input function data has different numbers of points and locations because the masked information varies among them.
Figure 18: 2D Nonlinear Darcy Example: The case with 75th percentile output function prediction error. The top left figure shows the queried input function data $u(\boldsymbol{x})$, with black circle symbols indicating the queried data. The middle top figure shows the input field reconstruction via the learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
Figure 19: 2D Nonlinear Darcy Example: The case with 50th percentile output function prediction error. The top left figure shows the queried input function data $u(\boldsymbol{x})$, with black circle symbols indicating the queried data. The middle top figure shows the input field reconstruction via the learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
Appendix C Burgers’ Equation
@@
Figure 21(a) illustrates the evolution of the loss function during the operator learning stage. The distribution of output function prediction errors after training is depicted in Figure 21(b). These results indicate that the training was successful, leading to accurate output function predictions by RINO.
@@ Appendix D Nonlinear 2D Darcy’s Equation on Unstructured Meshes
The evolution of the loss function during the operator learning stage is depicted in Figure 23(a) where both branch and trunk nets are unknown a priori. The distribution of output function prediction errors after training is plotted in Figure 23(b). A successful training with satisfactory accuracy in RINO predictions is confirmed by the results.
Figure 23: 2D Darcy’s Equation on Unstructured Meshes: (a) Loss function during optimization iterations for operator learning. (b) Distribution of output function prediction errors for the training and testing datasets after training.
As discussed in the main text, the proposed method can flexibly use a predefined trunk network by substituting a learned dictionary of approporiate basis functions for the output function data. Figure 24 shows the reconstruction error distributions for the output function point cloud data, where Algorithm 2 was able to identify an appropriate dictionary of 54 basis functions that span the output solution fields.
@@
The evolution of the loss function during the operator learning stage for the case where we utilized the learned, predefined trunk dictionary is depicted in Figure 25(a), where only the branch network is trained. The distribution of output function prediction errors after training is plotted in Figure 25(b). A successful training with satisfactory accuracy is confirmed by the results for the case where we indirectly learned a mapping between the coefficients of two sets of weakly orthogonal basis functions: one corresponding to the input function data and the other to the output function data, which were identified separately and offline.
Figure 25: 2D Darcy’s Equation on Unstructured Meshes: (a) Loss function during optimization iterations for operator learning with the predefined dictionary learned trunk net. (b) Distribution of output function prediction errors for the training and testing datasets after training.
@@ Appendix E Network Architecture
In this work, the neural network architectures and hyperparameters were chosen through manual trial and error and, therefore, are not optimized. As a result, the performance may improve by utilizing rigorous hyperparameter optimization methods, such as Bayesian optimization [38].
The hyperparaemtrs utilized for learning the dictionary are listed in Table LABEL:tab:hyper-dl. The column “2D Darcy*” corresponds to the last numerical example, where we solve the nonlinear 2D Darcy problem on different finite element meshes from one realization to another.
@@
Table 2: Hyperparameters for the dictionary learning algorithm where SIRENs are used as basis functions.
Antiderivative 1D Darcy 2D Darcy Burger’s 2D Darcy*
Number of hidden units $20$ $30$ $50$ $30$ $20$
Number of hidden layers $2$ $2$ $3$ $2$ $2$
Learning rate $1.66\times 10^{-4}$ $3\times 10^{-4}$ $6.6\times 10^{-4}$ $6.6\times 10^{-4}$ $4\times 10^{-4}$
$\lambda$ $10^{-4}$ $10^{-4}$ $10^{-5}$ $10^{-5}$ $10^{-3}$
$\omega_{0}$ $5$ $10$ $10$ $10$ $10$
@@
The same learning rate of 0.001 is chosen for the operator learning task. The input data to the branch and trunk networks are not normalized in all cases except for the nonlinear 2D Darcy’s equation, where the sparse codes obtained from the dictionary are linearly normalized to have a minimum of 0 and a maximum of 1 before training.
@@
In the case of the antiderivative, no branch network is utilized; the sparse codes obtained from the dictionary are directly considered as coefficients for the trunk basis, making the branch network essentially an identity operator. The architecture of the hidden layers in the branch networks for other examples are as follows:
  • Nonlinear 1D Darcy: MLP[50|ReLU|50]
  • Nonlinear 2D Darcy: MLP[50|ReLU|100]
  • Burgers’ equation: MLP[100|Tanh|100|Tanh|70]
  • Nonlinear 2D Darcy*: MLP[100|Tanh|100]
The architecture of the hidden layers in the trunk network are:
  • Antiderivative: SIREN[50,50] with $w_{0}=5$
  • Nonlinear 1D Darcy: SIREN[50,50,50] with $w_{0}=5$
  • Nonlinear 2D Darcy: MLP[50|ReLU|50|ReLU|100]
  • Nonlinear 2D Darcy*: SIREN[100,100,100] with $w_{0}=10$
For the “nonlinear 2D Darcy*” case, we also compare it with the scenario where our dictionary learning algorithm is applied to output function realizations to learn output basis functions offline and use them as a predefined trunk net. In this setup, each output basis was parametrized by SIREN[20|20] with $w_{0}=10$. The sparsity parameter $\lambda=10^{-4}$ and the learning rate is $10^{-3}$.
@@ Appendix F Ablation Study: Comparison with Gappy Proper Orthogonal Decomposition and Random Projection
In this example, we aim to empirically demonstrate that Algorithm 1 can effectively identify (weakly) orthogonal basis functions that accurately represent point clouds sampled from a random function. To enable a comparison with the GPOD method, we configure the problem setup to be compatible with it. Our method shares similarities with GPOD but offers greater flexibility: it can operate with point cloud data instead of regular grids and, more importantly, it finds continuous, fully differentiable basis functions rather than finite-dimensional basis vectors. Moreover, with GPOD, one cannot query an arbitrary sensor location that differs from the union of sensor locations used in the training data due to its discrete nature, while our approach allows queries at any location.
We purposely generate several realizations of random functions $u(x)$ that are inherently correlated by sharing a common low-dimensional structure. The $i$-th realization is given by:
@@
where the coefficients are assumed to be normally distributed, $\alpha_{l}^{(i)}\sim\mathcal{N}(0,1)$, and the basis functions are the 6th-order Legendre polynomial, $\psi_{1}(x)=P_{6}(x)$; the cosine function, $\psi_{2}(x)=\cos(3.3\pi x)$; and the absolute value function, $\psi_{3}(x)=2|x|-1$, as depicted in Figure 26(a). The random functions are generated similarly to PCE, where deterministic functions with random coefficients are used to model stochastic processes. Clearly, such random functions can be fully characterized by at most three modes, each corresponding to one of these functions. 200 realizations of this random function, where each realization is discretized at 100 uniform sensor locations, establish our “complete” training dataset, as plotted in a matrix format in Figure 26(b). In this figure, each row represents a realization, each column corresponds to a sensor location, and the colors at the $(i,j)$ location indicate the value of the $i$-th realization at the $j$-th sensor location. The “gappy” version of the data is created by randomly masking $R\%$ of each realization. For example, in Figure 26(c), the mask matrix for masking $90\%$ of the information is presented.
Note that the basis functions used in the generative function, Eq. (24), are purposely chosen to not be orthogonal (see Figure 28(a)) and exhibit a multiscale nature, as they have different length scales. It may be of interest to investigate whether our algorithm can find orthogonal basis functions, how similar these basis functions are to those found by GPOD, and how closely they resemble the non-orthogonal basis functions used for generating the data. It is known that the basis functions learned by PCA-based algorithms, such as GPOD, will be orthogonal.
(a) Manufactured bases
@@
Figure 26: (a) Basis functions used to generate random functions. (b) Complete training data set generated from 200 such functions, each sampled at 100 sensor locations. (c) Example of the mask matrix representing 90% masked training data for each function, where white (value 1) indicates masked entries.
In Figure 27, we have plotted the discovered bases from GPOD and those discovered by our method at various levels of masked data, ranging from $50\%$ to $95\%$. With sufficient unmasked data (e.g., $R=50\%,75\%$), the first two modes (orange and blue curves) from both GPOD and our algorithm closely resemble the cosine function and the Legendre function, which were initially weakly orthogonal in the generative process, as shown in Figure 28(a). However, the third mode (green curve) differs from the absolute value function, which was the non-orthogonal basis in the generative process, see Figure 28(a). From the GPOD perspective, this is expected since it represents data using orthogonal bases. Interestingly, the third basis function in our method exhibits almost the same behavior as the third basis in GPOD. Furthermore, the order of the bases in our approach matches the order in GPOD. These similarities occur even though we only weakly enforce the orthogonality constraint and do not enforce ordering based on data variance, as is the case with GPOD. The (weak) orthogonality of the bases in our method can be confirmed by the inner product calculation shown in Figure 28(c).
@@
Up to a certain level, GPOD results are consistent and can achieve almost zero reconstruction error. However, at higher levels of masking (e.g., $R=90\%$), its accuracy drops significantly. Our method, on the other hand, can still operate with reduced accuracy under similar conditions. We hypothesize that this is due to our method’s inductive bias towards smoothness in the basis functions, as they are parameterized by SIREN, whereas GPOD does not incorporate such an assumption. Nonetheless, our method also becomes less effective when the data is extremely sparse, as evidenced by the case with $R=95\%$.
@@
Figure 28: Orthogonality check among basis functions: (a) Basis functions used to generate random function data (see Eq. 24). (b) Basis vectors identified by the Gappy POD algorithm, and (c) basis functions identified by the proposed algorithm with 50% of the data masked.
The reconstruction errors at various level of masking percentages are listed in Table 3. The GPOD is a deterministic algorithm and, hence, there is no randomness regarding its prediction accuracy. However, our method relies on INRs as basis functions, where different randomly initialized parameters for the neural network may result in different predictions. Therefore, we conducted the training five times, each with different initial neural network parameters. As the results suggest, the accuracy of GPOD in this problem is better than Algorithm 1 and is resilient to the amount of masked data up to a certain limit, after which it suddenly fails catastrophically. On the other hand, our scheme gradually reduces accuracy with respect to the amount of masked data and remains performant under conditions where GPOD fails. This abrupt reduction in accuracy might be undesirable, particularly for data acquisition tasks where a well-behaved continuous error feedback is important.
$R\%$ Test Error (mean $\pm$ std) Train Error
@@
Table 3: Reconstruction errors for various randomly masked data. The same masking percentage $R$ is used for both training and testing.
@@
In Figure 29, we demonstrate the generalization ability of the proposed method. The shaded regions in green and purple represent two types of generalization: in-sample and out-of-sample in-distribution. If one attempts per-sample curve fitting, there is no guarantee that such fitting will capture the underlying ground truth, as fluctuations in these regions exceed the data frequency for interpolation; according to the Nyquist sampling theorem, the sampling rate should be at least twice the frequency of changes. Particularly under this scenario, the purple region can be considered pure extrapolation, which is known to be challenging without additional assumptions. The reason our reconstruction method performs so well is that it identifies common structures by considering all samples (realizations), not just individual samples, and utilizes basis functions that harness this structure. Moreover, although the available data in these samples are highly asymmetric, the reconstruction is almost symmetric. This symmetry is due to the identified symmetric basis functions, which are one of the main characteristics of the generated data, as the generative basis functions are symmetric. This property cannot be achieved through per-sample curve fitting. Therefore, we do not advocate for methods that attempt per-sample curve fitting and then align the samples by interpolation at fixed sensor locations.
@@
Figure 29: Reconstruction generalization capability of the proposed dictionary learning method, demonstrated on two queried samples, with the dictionary trained on only 15% of the data.
@@
One may consider using random smooth functions, such as those parameterized by a shallow neural network [33, 31, 32, 16], as a method to construct basis functions in a predefined manner. These functions can naturally impose smoothness as inductive biases and can remain resolution-independent during the reconstruction and projection steps. Here we aim to benchmark and compare their performance using this pedagogical example. To this end, we consider two cases: 1) random shallow ReLU networks, where each basis is a ReLU network with 20 hidden neurons randomly initialized, and 2) classical cosine basis functions with random frequency and phase shifts as follows:
@@
for the $i$-th cosine basis functions. For completeness, we also consider the classical monomial and Legendre basis functions (deterministic). In Figure 30, we compare the reconstruction performance of these sets of basis functions by increasing the number of basis functions used in each set and by gradually sparsifying the data.
When the data density is sufficiently high, as shown in Figure 30(a), random ReLU, random cosine, and Legendre bases deliver satisfactory performance. However, in our numerical experiments, they appear to require two orders of magnitude more basis functions than ours, which is not ideal for downstream learning tasks, as one would prefer to operate with an embedding space of dimension 3 rather than 100. This preference is due to both computational efficiency and the fact that the amount of data required for such learning tasks may scale exponentially with respect to dimensionality. As data sparsity increases, as shown in Figure 30(b and c), the performance gap between our method and others grows more pronounced. For instance, with 85% of the data masked in Figure 30(c), the performance of the other methods becomes unacceptable. This decline is due to the limited data, making these other methods more prone to overfitting.
@@
Figure 30: Comparison of reconstruction accuracy on test data using four distinct sets of basis functions. Each set is evaluated through five independent computations, with randomly chosen parameters for stochastic methods; for deterministic basis functions (monomial and Legendre), no randomness is involved. As the results demonstrate, even as data sparsity increases, our method consistently outperforms alternatives. Notably, our approach achieves this with significantly fewer basis functions, utilizing only 3 compared to 100 in other methods.
To further analyze the performance of random bases in the case where the data has the highest sparsity (i.e., 85% masked data), we plotted the reconstruction curves for the worst case and the 25th quintile in terms of relative errors in Figure 31. As expected (and as consistent with the literature [31, 16]), the random basis functions perform well in the interpolation regime but poorly in the extrapolation regime beyond the data support (as observed in the purple-shaded areas in these figures). Moreover, the underlying generative process of the data is symmetric at $x=0$, see Figure 26(a). However, there is no guarantee that this symmetry can be recovered using random basis functions. As already discussed, these issues do not arise with our method, as we learn appropriate basis functions by considering the common correlation structure across all realizations, rather than merely interpolating for each individual realization.
@@
Figure 31: Reconstruction curves for the worst case and the 25th quantile in terms of relative reconstruction error in the case where 85% of the data was masked.

### v3

@@ 5 Acknowledgements
We acknowledge the discussion with Professor George E. Karniadakis during the CRUNCH seminar, who brought random projection and ReNO [3] to our attention, which improved the work.
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number DE-SC0024162.
@@
Supplementary results for the numerical examples are provided in Appendices A, B and C. Appendix E lists the hyperparameters used for each problem. An ablation study for Algorithm 1, comparing the RINO method with GPOD, is presented in Appendix F. Moreover, the use of random projections as an alternative for finding vector representations of input or output function data has been studied and compared to the proposed dictionary learning algorithm. Additional potential advantages of RINO over RI-DeepONet in terms of generalization are examined in Appendix G. Appendix H summarizes the formulation for Gaussian random field generation.
@@ Appendix A Nonlinear 1D Darcy’s Equation
The history of the loss function during the operator learning stage is plotted in Figure 20(a). In Figure 20(b), the distribution of output function prediction errors after training is shown. The results suggest that the training is successful and the output function prediction via RI-DeepONet is very accurate.
@@ Appendix B Nonlinear 2D Darcy’s Equation
Two examples of the test predictions for the output field are shown in Figures 23 and 24, corresponding to the 75th percentile error and the median error, respectively. In each example, the queried point cloud input function data has different numbers of points and locations because the masked information varies among them.
Figure 23: 2D Nonlinear Darcy Example: The case with 75th percentile output function prediction error. The top left figure shows the queried input function data $u(\boldsymbol{x})$, with black circle symbols indicating the queried data. The middle top figure shows the input field reconstruction via the learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the output field $s(\boldsymbol{y})$, the RI-DeepONet prediction, and its corresponding errors.
Figure 24: 2D Nonlinear Darcy Example: The case with 50th percentile output function prediction error. The top left figure shows the queried input function data $u(\boldsymbol{x})$, with black circle symbols indicating the queried data. The middle top figure shows the input field reconstruction via the learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the output field $s(\boldsymbol{y})$, the RI-DeepONet prediction, and its corresponding errors.
Appendix C Nonlinear Burgers’ Equation
@@
Figure 26(a) illustrates the evolution of the loss function during the operator learning stage. The distribution of output function prediction errors after training is depicted in Figure 26(b). These results indicate that the training was successful, leading to accurate output function predictions by RI-DeepONet.
@@ Appendix D Nonlinear 2D Darcy’s Equation on Unstructured Meshes
As discussed in the main text, the proposed RINO method can build the neural operator by learning approporiate basis functions for the output function data as well. Figure 28 shows the reconstruction error distributions for the output function point cloud data, where Algorithm 2 was able to identify an appropriate dictionary of 37 basis functions that span the output solution fields.
@@ Appendix E Network Architecture
In this work, the neural network architectures and hyperparameters were chosen through manual trial and error and, therefore, are not optimized. As a result, the performance may improve by utilizing rigorous hyperparameter optimization methods, such as Bayesian optimization [55]. In the following, “2D Darcy*” corresponds to the numerical example where we solve the nonlinear 2D Darcy problem on different finite element meshes from one realization to another.
E.1 Input basis functions
@@
The hyperparameters utilized for learning the dictionary of input basis functions are listed in Table LABEL:tab:hyper-dl-in.
Table 2: Hyperparameters of SIREN input basis functions used in the dictionary learning algorithm
Antiderivative 1D Darcy 2D Darcy Burger’s 2D Darcy* 2D Elasticity
Number of hidden units $20$ $30$ $50$ $30$ $20$ $100$
Number of hidden layers $2$ $2$ $3$ $2$ $2$ $2$
Learning rate $1.66\times 10^{-4}$ $3\times 10^{-4}$ $6.6\times 10^{-4}$ $6.6\times 10^{-4}$ $4\times 10^{-4}$ $5\times 10^{-4}$
$\lambda$ $10^{-4}$ $10^{-4}$ $10^{-5}$ $10^{-5}$ $10^{-3}$ $10^{-5}$
$\omega_{0}$ $5$ $10$ $10$ $10$ $10$ 20
E.2 Output basis functions
The hyperparameters utilized for learning the dictionary of output basis functions are listed in Table LABEL:tab:hyper-dl-out. In the case of 2D solid mechanics, $\omega_{0}$ is set to 35, 25, and 15 for stress components $s_{11}$, $s_{12}$, and $s_{22}$, respectively.
Table 3: Hyperparameters of SIREN output basis functions used in the dictionary learning algorithm
2D Darcy* 2D Solid Mechanics
Number of hidden units $20$ $100$
Number of hidden layers $2$ $2$
Learning rate $1\times 10^{-3}$ $5\times 10^{-4}$
$\lambda$ $10^{-4}$ $10^{-5}$
$\omega_{0}$ $10$ 35, 25, 15
E.3 Operator learning
@@
The obtained embeddings are not normalized in any case except for the nonlinear 2D Darcy equation, where the sparse codes from the dictionary are linearly normalized to range between 0 and 1 prior to training. The learning rate of 0.001 is chosen for all operator learning tasks.
@@
Antiderivative: RI-DeepONet is employed without a branch network; the sparse codes obtained from the dictionary are directly used as coefficients for the trunk basis, effectively making the branch network an identity operator. The trunk network is SIREN with two hidden layers of size 50 and $w_{0}=5$.
Nonlinear 1D Darcy: RI-DeepONet is employed with a branch network consisting of an MLP with two hidden layers, each of size 50, and ReLU activation, and a trunk network implemented as a SIREN with three hidden layers, each of size 50, and a frequency parameter $w_{0}=5$.
Nonlinear 2D Darcy: RI-DeepONet is employed with a branch network consisting of an MLP with two hidden layers of sizes 50 and 100, using ReLU activation, and a trunk network implemented as an MLP with three hidden layers of sizes 50, 50, and 100, also using ReLU activation.
Nonlinear Burgers’ equation: RI-POD-DeepONet is employed with a branch network consisting of three hidden layers of sizes 100, 100, and 70 using Tanh activation.
Nonlinear 2D Darcy*: RINO is employed, where the network $\mathcal{F}^{\text{ri}}$ is an MLP with a single hidden layer of size 40, using Tanh activation. The input and output embeddings have dimensionality 40 and 37, respectively.
Solid Mechanics: RINO is employed, where the networks $\mathcal{F}^{\text{ri}}$ are MLPs with three hidden layers and Tanh activation functions. The number of hidden units are 100, 50, and 50 for stress components $s_{11}$, $s_{12}$, and $s_{22}$ respectively. These MLPs are used to map input embedding of dimension 24 to output embeddings of dimensions 41, 29, and 14 for $s_{11}$, $s_{12}$, and $s_{22}$ respectively.
@@ Appendix F Ablation Study: Comparison with Gappy Proper Orthogonal Decomposition and Random Projection
In this example, we aim to empirically demonstrate that Algorithm 1 can effectively identify continuous (weakly) orthogonal basis functions that accurately represent point clouds sampled from a random function. We compare the proposed dictionary learning to GPOD, but we note first some important distinctions. The proposed method shares certain similarities with GPOD, but offers greater flexibility. In particular, it can operate with point cloud data instead of regular grids and, more importantly, it finds continuous, fully differentiable basis functions rather than finite-dimensional basis vectors. Moreover, with GPOD, one cannot query an arbitrary sensor location that differs from the union of sensor locations used in the training data due to its discrete nature, while the proposed approach allows queries at any location.
For the purpose of comparison, we generate several realizations of random functions $u(x)$ that share a common low-dimensional structure defined through the following expansion where the $i$-th realization is given by:
@@
the coefficients are assumed to be normally distributed, $\alpha_{l}^{(i)}\sim\mathcal{N}(0,1)$, and the arbitrary basis functions are the 6th-order Legendre polynomial, $\psi_{1}(x)=P_{6}(x)$; the cosine function, $\psi_{2}(x)=\cos(3.3\pi x)$; and the absolute value function, $\psi_{3}(x)=2|x|-1$, as depicted in Figure 29(a). The random functions are generated similarly to PCE, where deterministic functions with random coefficients are used to model stochastic processes. Clearly, such random functions can be fully characterized by at most three modes, each corresponding to one of these functions. We generate 200 realizations of this random function, where each realization is discretized at 100 uniform sensor locations to establish the “complete” training dataset, as plotted in the matrix in Figure 29(b). In this figure, each row represents a realization, each column corresponds to a sensor location, and the colors at the $(i,j)$ location indicate the value of the $i$-th realization at the $j$-th sensor location. The “gappy” version of the data is created by randomly masking $R\%$ of each realization. For example, in Figure 29(c), the mask matrix for masking $90\%$ of the information is presented.
(a) Arbitrary bases
@@
Figure 29: (a) Arbitrary basis functions used to generate random functions. (b) Complete training data set of 200 random functions generated from the arbitrary basis in (a), each sampled at 100 sensor locations. (c) Example of the mask matrix representing 90% masked training data for each function, where white (value 1) indicates masked entries.
Note that the arbitrary basis functions used to generate the random functions from Eq. (25), are purposely chosen to not be orthogonal (see Figure 31(a)) and exhibit a multiscale nature, as they have different length scales. We seek to investigate whether the proposed dictionary learning algorithm can find orthogonal basis functions, how similar these basis functions are to those found by GPOD, and how closely they resemble the non-orthogonal basis functions used to generate the data. It is well-known that the basis vectors learned by PCA-based algorithms, such as GPOD, will be orthogonal.
In Figure 30, the discovered bases from GPOD and the proposed dictionary learning method are plotted for various levels of masked data, ranging from $50\%$ to $95\%$. With sufficient unmasked data (e.g., $R=50\%,75\%$), the first two modes (orange and blue curves) from both GPOD and dictionary learning closely resemble the cosine function and the Legendre function, which were initially weakly orthogonal in the generative process, as shown in Figure 31(a). However, the third mode (green curve) differs from the absolute value function, which was the non-orthogonal basis in the generative process, see Figure 31(a). From the GPOD perspective, this is expected since it represents data using orthogonal bases. Interestingly, the third basis function from the proposed dictionary learning closely resembles the third basis from GPOD. Furthermore, the order of the bases in the proposed dictionary learning approach matches the order in GPOD. These similarities occur even though we only weakly enforce the orthogonality constraint and do not enforce ordering based on data variance, as is the case with GPOD. The (weak) orthogonality of the bases in the proposed method can be confirmed by the inner product calculation shown in Figure 31(c).
@@
Figure 30 further shows that, up to a certain level of data masking (75%), the GPOD results are consistent with dictionary learning and Table 4 shows that it can achieve very small reconstruction error. However, at higher levels of masking (e.g., $R\geq 90\%$), its accuracy drops significantly and the basis vectors become noisy. The proposed dictionary learning method, on the other hand, still produces smooth, orthogonal basis functions with low reconstruction error. In fact, even in the extreme case of $R=95\%$, the basis functions remain smooth and the accuracy, though, diminished is much better than GPOD, which has extremely noisy basis vectors.
@@
Figure 31: Orthogonality check among basis functions: (a) Basis functions used to generate the random function data (see Eq. 25); (b) Basis vectors identified by the Gappy POD algorithm with 50% of the data masked; and (c) Basis functions identified by the proposed dictionary learning algorithm with 50% of the data masked.
The reconstruction errors at various level of masking percentages are listed in Table 4 for both GPOD and the proposed dictionary learning. Note that the GPOD is a deterministic algorithm and, hence, there is no variability in its prediction accuracy. However, the proposed dictionary learning algorithm relies on INRs as basis functions, where different randomly initialized parameters for the neural network result in slightly different predictions. Therefore, we repeated the training five times, each with different initial neural network parameters. As the results suggest, for small masking percentage the accuracy of GPOD in this problem is better than Algorithm 1 and is resilient to the amount of masked data up to a certain limit, after which it suddenly fails catastrophically. On the other hand, the proposed scheme is still very accurate for small masking percentage and gradually reduces accuracy as the amount of masked data increases. Importantly, it continues to perform well under high masking rates where GPOD fails. This abrupt reduction in accuracy might be undesirable, particularly for data acquisition tasks where a well-behaved continuous error feedback is important.
$R\%$ Train Error Test Error (mean $\pm$ std)
@@
Table 4: Reconstruction errors for various percentages of randomly masked data. The same masking percentage $R$ is used for both training and testing.
@@
The most important feature of the proposed dictionary learning is that we learn smooth, continuous orthogonal basis functions from all realizations, which improves generalization in predictions. In Figure 32, we demonstrate the generalization ability of the proposed method. The shaded regions in green and purple represent two types of generalization: interpolation and extrapolation of in-distribution functions. If one attempts per-sample curve fitting, there is no guarantee that such fitting will capture important features of the underlying ground truth function, as fluctuations in these regions exceed the data frequency for interpolation; according to the Nyquist sampling theorem, the sampling rate should be at least twice the frequency of changes. Particularly under this scenario, the purple region can be considered pure extrapolation, which is known to be challenging without additional assumptions. The reason the proposed reconstruction method performs so well is that it identifies common structures in the data set by considering all samples (realizations), not just individual samples, and utilizes basis functions that harness this structure. Moreover, although the available data in these samples are highly asymmetric, the reconstruction is almost symmetric. This symmetry is due to the identified symmetric basis functions, which are one of the main characteristics of the generated data, as the generative basis functions are symmetric. This property cannot be achieved through per-sample curve fitting. Therefore, we do not advocate for methods that attempt per-sample curve fitting and then align the samples by interpolation at fixed sensor locations.
@@
Figure 32: Reconstruction generalization capability of the proposed dictionary learning method, demonstrated on two queried samples, with the dictionary trained on only 15% of the data. Purple areas highlight regions of extrapolation and green areas show regions where the model interpolates well despite lacking data elucidating features in these regions.
Given that the advantages of the proposed dictionary learning method derive from the continuous nature of the basis functions, one may consider assigning the basis arbitrarily. For example, we’ve already mentioned the possiblity of using orthogonal polynomials. Another options is to define the basis using random smooth functions, such as those parameterized by a shallow neural network [48, 46, 47, 19] – a method referred to as random projection. These functions can naturally impose smoothness as inductive biases and can remain resolution-independent during the reconstruction and projection steps.
@@
Here we aim to benchmark and compare the performance of the proposed dictionary learning method to different methods that employ arbitrary pre-defined continuous basis functions using the pedagogical example described above (Eq. (25)). To this end, we consider four cases: 1) random shallow ReLU networks, where each basis is a ReLU network with 20 hidden neurons randomly initialized; 2) classical cosine basis functions with random frequency and phase shifts as follows:
@@
for the $i$-th cosine basis functions; 3) the classical monomial basis functions; and 4) Legendre polynomial basis functions (deterministic). In Figure 33, we compare the reconstruction accuracy using these different sets of basis functions for an increasing number of basis functions and different degrees of random data masking.
@@
Figure 33: Comparison of reconstruction accuracy on test data using four distinct sets of basis functions: Random projections with shallow ReLU networks (orange), Random projections with cosine functions (green), Monomials (red), Legendre Polynomials (blue), and the proposed dictionary learning (purple). Each set is evaluated for five independent trials with randomly chosen parameters for stochastic methods. For deterministic basis functions (monomial and Legendre), no randomness is involved and a single trial is performed. As the results demonstrate, even as data sparsity increases, the proposed dictionary learning method consistently outperforms alternatives. Notably, dictionary learning achieves lower reconstruction error with significantly fewer basis functions, needing only 3 basis functions compared to hundreds in other methods.
When the data density is sufficiently high, as shown in Figure 33(a), the random ReLU, random cosine, and Legendre bases deliver satisfactory performance. However, they appear to require two orders of magnitude more basis functions than the proposed dictionary learning method, which presents challenges for downstream learning tasks, as operations in embedding space of dimension 3 are much simpler than operations in an embedding space of 100 or more dimensions. This is due to both computational efficiency and the fact that the amount of data required for such learning tasks may scale exponentially with dimensionality. As data sparsity increases, as shown in Figure 33(b and c), the performance gap between the proposed dictionary learning method and others grows more pronounced. For instance, with 85% of the data masked in Figure 33(c), the performance of the other methods becomes unacceptable. This decline is due to the limited data, making these other methods prone to overfitting.
To further analyze the performance of random bases in the case where the data has the highest sparsity (i.e., 85% masked data), we plotted the reconstruction curves for the worst case and the 25th quintile in terms of relative errors in Figure 34. As expected (and as consistent with the literature [46, 19]), the random basis functions perform well in the interpolation regime but poorly in the extrapolation regime beyond the data support (as observed in the purple-shaded areas in these figures). Moreover, the underlying generative process of the data is symmetric at $x=0$, see Figure 29(a). However, there is no guarantee that this symmetry can be recovered using random basis functions. As already discussed, these issues do not arise in the proposed dictionary learning method, as we learn appropriate basis functions by considering the common correlation structure across all realizations, rather than merely interpolating for each individual realization.
@@
Figure 34: Reconstruction curves for the worst case (left pane) and the 25th quantile (right pane) in terms of relative reconstruction error in the case where 85% of the data was masked.